OpenStudy (anonymous):

what is -1 4/5 x (3 1/3)

OpenStudy (studygurl14):

\[-1\frac{4}{5} \times 3 \frac{1}{3}\] Convert \[-\frac{9}{5} \times \frac{10}{3}\] What now?